# Tony Lindeberg

### v3

@@ Abstract
We propose that this notion of time-causal wavelet analysis could be a valuable tool for signal processing tasks, where streams of signals are to be processed in real time, specifically for signals that may contain local variations over a rich span of temporal scales, or more generally for analysing physical or biophysical temporal phenomena, where a fully time-causal analysis is called for to be physically realistic.
@@ 1 Introduction
Temporal signals can inherently contain different types of structures at different temporal scales. For the purpose of handling such scale variations in, not necessarily, only temporal signals, the notion of wavelet analysis has been developed by the signal processing community, to decompose signals into different components at different scales (Ricker 1944, Morlet et al. 1982a; 1982b, Grossmann and Morlet 1984, Hosken 1988, Mallat 1989; 1999, Heil and Walnut 1989, Strang 1989, Daubechies 1992, Meyer 1992, Chui 1992, Ruskain et al. 1992, Donoho and Johnston 1993, Teolis 1998, Walker 1999, Debnath and Shah 2002, Misiti et al. 2007).
@@
The subject of this paper is to show how a temporal wavelet theory can be developed based on the notion of time-causal temporal scale-space representation, originally developed for video analysis in Lindeberg (2016) and then dedicated for a pure time-causal temporal domain in Lindeberg (2023). This temporal scale-space theory is based on convolution with a special temporal smoothing kernel, denoted the time-causal limit kernel, and which corresponds to the convolution of an infinite set of truncated exponential kernels in cascade, with especially chosen time constants to obtain temporal scale covariance.
@@
By the time-causal limit kernel constituting the composition of a set of truncated exponential kernels in cascade, the temporal smoothing operation equivalently corresponds to applying a set of first-order integrators in cascade. In terms of discrete implementation, such an operation can be approximated by a set of first-order recursive filters coupled in cascade. Thereby, the resulting implementation is fully time-recursive, meaning that no other temporal memory of the past is needed, beyond the information contained in the temporal scale channels themselves. Specifically, this implies straightforward implementation schemes on regular signal processing architectures.
@@
In this way, we will demonstrate that significant components in more traditional non-causal wavelet theory can be carried over to a time-causal temporal domain, which opens up for a variety of application areas, where either (i) temporal signals are to be processed in real time, such as for purposes of real-time monitoring or as components in control systems, that are to react to a continuous inflow of information, or (ii) analysis of biological or physical signals that needs to be performed in a completely time-causal manner, to be biologically or physically realistic, such as in mathematically based models of biological or physical processes.
@@ 1.1 Structure of the presentation
Section 4 then shows how this continuous kernel can be canonically discretized in terms of a set of first-order recursive filters coupled in cascade, complemented with time-causal difference operators to compute discrete approximations of temporal derivatives of the time-causal limit kernel, thus leading to a family of discrete time-causal wavelet representations as summarized in Section 5.
@@
Finally, Section 8 concludes with a summary and discussion, including relations to related multi-scale signal processing approaches.
@@ 2 Continuous wavelet representations based on temporal derivatives of the time-causal limit kernel
In this section, we will after (i) a first review of main concepts regarding regular, not necessarily time-causal, wavelet theory, describe (ii) the time-causal limit kernel and how this kernel is constrained with respect its shape based on (iii) a very general theoretical result regarding provably variation-diminishing convolution transformations and the desirable property of basing the time-causal wavelet analysis on a (iv) provably scale covariant convolution kernel.
@@
Based on these foundations, we will then describe (viii) how the first- and second-order temporal derivatives of the time-causal limit kernel can be regarded as a time-causal analogues of the Ricker wavelet, also referred to as the Mexican hat wavelet, and corresponding to the second-order derivative of the Gaussian kernel, as well as analogues of the first-order derivatives of the non-causal Gaussian kernel.
@@ 2.2 The time-causal limit kernel
over multiple temporal scale levels $\tau\in{\mathbb{R}}_{+}$ in such a way that no new local extrema or zero-crossings may be created from finer to coarser levels of temporal scales. The parameter $c$ is specifically a distribution parameter with $c>1$ that determines the discrete scale levels in units of the variance of the time-causal limit kernel according to [^2] This logarithmic spacing of the temporal scale levels is similar to the spacing generated by a uniform spacing with respect to the canonical logarithmic parameterization of the scale parameter into an effective scale parameter $\tau_{\text{eff}}$ according to $\tau_{\text{eff}}=A+B\log\tau$, as can be derived from axiomatic arguments for a regular spatial scale-space representation (Lindeberg 1993). Compared to a uniform spacing of the scale levels relative to the scale parameter measured in units of the variance $\tau$ of the temporal smoothing kernel, the logarithmic distribution according to (12) also leads to significantly shorter delays, as further described in Lindeberg (2016) Section 4 and Appendix 1.
@@
where $\tau_{0}>0$ represents a reference scale, and where for common purposes of implementation one often chooses either $c=\sqrt{2}$ or $c=2$ and $\tau_{0}=1$. Alternatively, the time-causal limit kernel can also be parameterized in terms of the standard deviation $\sigma=\sqrt{\tau}$ of the time-causal limit kernel, which is of dimension $\mbox{time}]$.
@@ 2.4 Scale covariance property of the time-causal limit kernel
Temporal scale covariance here means that if we for any temporal scaling factor $S_{t}$ that is an integer power of the distribution parameter $c>1$ according to
$S_{t}=c^{j}\quad\quad\mbox{for}\quad\quad j\in{\mathbb{Z}}$ (20)
@@
This result thus means that the time-causal limit kernel is self-similar under temporal scaling transformations, which also, with appropriate temporal scale normalization, carries over to temporal derivatives of the time-causal limit kernel according to
@@
Note, however, that since the temporal scale levels are inherently discrete, as implied by the construction of the temporal scale-space representation from multiple truncated exponential kernels coupled in cascade, the temporal scale covariance property holds exactly only when the temporal scaling factor $S_{t}$ is an integer power of the distribution parameter $c$. For other values of $S_{t}$, the corresponding result will only be approximate.
@@ 2.5.3 Trade-off issue regarding the choice of the distribution parameter $c$
  • A larger of the distribution parameter $c$ will lead to shorter temporal delays, as estimated according to (28) and (29), with better ability to respond fast in real-time situations.
@@ 2.5.4 Differential equation formulation of the time-causal temporal scale-space representation
In terms of differential equations, [^4] The fact that the time-causal scale-space representation $L(t;\;\tau,c)$ according to (11) satisfies the recurrence relation (30) between adjacent levels of temporal scales can be formally verified as follows: Let us first rewrite the recurrence relation between the input signal $f_{\text{in}}(t)$ and the output signal $f_{\text{out}}(t)$ between adjacent levels of temporal scales $\tau_{k-1}$ and $\tau_{k}$ as $\partial_{t}f_{\text{out}}(t)=-1/\mu_{k}\times(f_{\text{out}}(t)-f_{\text{in}}(t))$. Taking the Laplace transform of this relation then gives $s\,{\cal L}(f_{\text{out}})(s)=-1/\mu_{k}\times({\cal L}(f_{\text{out}})(s)-{\cal L}(f_{\text{in}})(s))$. Let us next assume that the Laplace transforms ${\cal L}(f_{\text{out}})(s)$ and ${\cal L}(f_{\text{in}})(s)$ of $f_{\text{out}}(t)$ and $f_{\text{in}}(t)$, respectively, are related by a transfer function $H(s)$ according to ${\cal L}(f_{\text{out}})(s)=H(s)\,{\cal L}(f_{\text{in}})(s)$. Inserting this expression into the Laplace transform of the recurrence relation, then gives $s\,H(s)\,{\cal L}(f_{\text{in}})(s)=-1/\mu_{k}\times(H(s)\,{\cal L}(f_{\text{in}})(s)-{\cal L}(f_{\text{in}})(s))$, from which we can solve for the transfer function as $H(s)=1/(1+\mu_{k}\,s)$. From tables of the Laplace transforms of standard functions, we have that the Laplace transform of the truncated exponential kernel $e^{-\alpha\,t}$ for $t>0$, 0 elsewhere, is $1/(s+\alpha)$. From this, we can conclude that the transfer function $H(s)$ corresponds to the Laplace transform of the truncated exponential kernel $h_{\exp}(t;\;\mu_{k})=1/\mu_{k}\times e^{-t/\mu_{k}}$ for $t>0$, 0 elsewhere, which proves the result. the relationships between adjacent levels of temporal scales are given by
@@
An important property of this type of temporal scale-space representation is that it is time-recursive. The temporal scale-space representations $L(t;\;\tau_{k},c)$ over the temporal scale channels with indices $k$ constitute a sufficient temporal memory of the past to compute the temporal scale-space representation and the next temporal moment, given a new input in the input signal $f(t)$.
The form of the first-order integrators (30) also explicitly implies that local perturbations in the signal are gradually smoothed out from finer to coarser levels of temporal scales, because of the negative feedback with respect to the differences $L(t;\;\tau_{k},c)-L(t;\;\tau_{k-1},c)$ between adjacent temporal scale channels.
@@ 2.5.5 Cascade smoothing property of the time-causal limit kernel
This property means that, when computing the temporal scale-space representation or a temporal derivative of the temporal scale-space representation, it is not necessary to redo all the computations when to compute those representations over multiple temporal scales. Instead, it is sufficient to perform the major amount of temporal smoothing once and for all, and then compute the representations at the coarser levels of temporal scales by applying a usually much smaller number of truncated exponential kernels to the corresponding representations at finer temporal scales.
@@ 2.6 Time-causal and time-recursive mother wavelets and wavelet representations
According to traditional wavelet theory, one usually makes use of $L_{2}$-norms when normalizing the mother wavelet. For purposes of computing derivatives over multiple scales in scale-space theory, however, other choices of $L_{p}$-norms are more common for relating image structures over scales (Lindeberg 1998 Table 3), as will be detailed further below.
@@ 2.7 Relations to scale-normalized derivatives in scale-space theory
For purposes in scale-space theory, the notion of scale-normalized temporal derivatives can be introduced according to (Lindeberg 1998 Equation (6); Lindeberg 2017 Equation (6))
@@ 2.7.1 Scale covariance property of scale-normalized temporal derivatives
This means that we can readily compare the magnitudes of scale-normalized derivatives between different temporal moments $\{t\leftrightarrow t^{\prime}\}$ and temporal scales $\{\tau\leftrightarrow\tau^{\prime}\}$.
@@ 2.7.2 Scale selection properties for characteristic temporal signals
Notably, these results bear close relationships to the matched filter theorem (Woodward 1953, Turin 1960), in that the scale selection mechanism will choose filters for detecting the different types of temporal structures in the input signal, that as well as possible match their size.
@@
In relation to normalization of the filters with respect to $L_{p}$-norms, it can specifically be shown that (Lindeberg 1998 Section 9.1) the notion of $\gamma$-normalized derivatives corresponds to normalizing the $n$:th order scale-normalized Gaussian derivative kernels $g_{\zeta^{n}}(t;\;\tau)$ over a one-dimensional domain to constant $L_{p}$-norms over scale
@@ 3.1 Time-causal wavelet representations based on temporal derivatives of the time-causal limit kernel
for the temporal scale levels chosen as
@@
In the following analysis to be performed, we will, however, not explicitly normalized the mother wavelets to constant $L_{p}$-norm, instead fully relying on the up to a uniform scaling factor normalization over scales in terms of scale-normalized temporal derivatives according to (45).

### v6

@@ Abstract
We propose that this notion of time-causal wavelet analysis could be a valuable multi-purpose tool for signal processing tasks, where streams of time-dependent signals are to be processed in real time, specifically for signals that may contain local variations over a rich span of temporal scales, or more generally for analysing physical or biophysical temporal phenomena, where a fully time-causal analysis is called for to be physically realistic.
@@ 1 Introduction
Temporal signals can inherently contain different types of structures at different temporal scales. Making such multi-scale relationships in the data explicit, and exploring these relationships, can often substantially improve the performance of signal processing pipelines. For the purpose of handling such scale variations in, not necessarily, only temporal signals, the notion of wavelet analysis has therefore been developed by the signal processing community, to decompose signals into different components at different scales (Ricker 1944, Morlet et al. 1982a; 1982b, Grossmann and Morlet 1984, Hosken 1988, Mallat 1989; 1999, Heil and Walnut 1989, Strang 1989, Daubechies 1992, Meyer 1992, Chui 1992, Ruskain et al. 1992, Donoho and Johnston 1993, Teolis 1998, Walker 1999, Debnath and Shah 2002, Misiti et al. 2007).
@@
The subject of this paper is to show how a temporal wavelet theory can be developed based on the notion of time-causal temporal scale-space representation, originally developed for video analysis in Lindeberg (2016), and then dedicated for a pure time-causal temporal domain in Lindeberg (2023). This temporal scale-space theory is based on convolution with a special temporal smoothing kernel, denoted the time-causal limit kernel, and which corresponds to the convolution of an infinite set of truncated exponential kernels in cascade, with especially chosen time constants to obtain temporal scale covariance.
@@
Furthermore, by forming the differences between temporal scale-space representations at adjacent temporal scales, we can construct time-causal bandpass wavelet representations, which are formally shown to up to a scaling factor correspond to the time-causal wavelet representation based on first-order temporal derivatives of the time-causal limit kernel. From such a temporal bandpass representation, reconstruction of the original signal is straightforward, by mere addition of the time-causal bandpass representations, which opens up possibilities for real-time manipulation of the signal as decomposed into a time-causal wavelet representation. This situation is structurally very closely related to the corresponding property of the Ricker or Mexican hat wavelet, based on second-order derivatives of the non-causal Gaussian kernel, where reconstruction can be directly performed by integrating the corresponding second-order non-causal wavelet representations over the scale parameter, or the difference-of-Gaussians wavelet, where reconstruction is performed by instead summing the difference-of-Gaussians wavelet representations over temporal scales.
By the time-causal limit kernel constituting the composition of a set of truncated exponential kernels in cascade, the temporal smoothing operation equivalently corresponds to applying a set of first-order integrators in cascade. In terms of discrete implementation, such an operation can be approximated by a set of first-order recursive filters coupled in cascade. Thereby, the resulting implementation is fully time-recursive, meaning that no other temporal memory of the past is needed, beyond the information contained in the temporal scale channels themselves. Specifically, this implies straightforward and very efficient implementation schemes on regular signal processing architectures.
@@
In this way, we will demonstrate that significant components in more traditional non-causal wavelet theory can be carried over to a time-causal temporal domain, which opens up for a variety of application areas, where either (i) temporal signals are to be processed in real time, such as for purposes of real-time monitoring, time series analysis and prediction, or as components in control systems, that are to react to a continuous inflow of information, or (ii) analysis of biological or physical signals that needs to be performed in a completely time-causal manner, to be biologically or physically realistic, such as in mathematically based models of biological or physical processes.
In this paper, we develop the underlying fundaments of the proposed class of time-causal wavelets, with special emphasis on their theoretical properties, combined with characterizations of scaling properties and proofs-of-concepts for different types of temporal signals.
Beyond a very few early approaches, this is the first systematic treatment of time-causal wavelet theory, with structurally close relations to non-causal wavelets based on derivatives of the Gaussian kernel. Specifically, we propose that the time-causal wavelet theory to be presented here constitutes a canonical way to define wavelet representations over a time-causal temporal domain, in a corresponding way as non-causal wavelets based on derivatives of the Gaussian kernel, such as the Ricker or Mexican hat wavelets, can be regarded as canonical wavelet representations over a non-causal temporal domain, according to the theory for variation-diminishing convolution transformations.
@@ 1.1 Structure of the presentation
Section 4 then shows how this continuous kernel can be canonically discretized in terms of a set of first-order recursive filters coupled in cascade, complemented with time-causal difference operators to compute discrete approximations of temporal derivatives of the time-causal limit kernel, thereby leading to a family of discrete time-causal wavelet representations as summarized in Section 5.
@@
Section 7 complements with examples of results of computing the proposed time-causal wavelet representations for different types of temporal signals. Finally, Section 8 concludes with a summary and discussion, including relations to related multi-scale signal processing approaches.
@@ 2 Continuous wavelet representations based on temporal derivatives of the time-causal limit kernel
In this section, we will after (i) a first review of main concepts regarding regular, not necessarily time-causal, wavelet theory, describe (ii) the time-causal limit kernel and how this kernel is constrained with respect its shape based on (iii) a very general theoretical result regarding provably variation-diminishing convolution transformations and the desirable property of basing the time-causal wavelet analysis on a (iv) provably scale-covariant convolution kernel.
@@
Based on these foundations, we will then describe (viii) how the first- and second-order temporal derivatives of the time-causal limit kernel can be regarded as a time-causal analogues of the Ricker wavelet, also referred to as the Mexican hat wavelet, and corresponding to the second-order derivative of the Gaussian kernel, as well as also analogues of the first-order derivatives of the non-causal Gaussian kernel.
Specifically, since the time-causal limit kernel, that we will use as a fundamental basis for defining the time-causal wavelet representations, does not have any compact closed-form expression, we will in this section devote special emphasis on describing its theoretical properties, so as to be able to infer properties of the derived time-causal wavelet representations, and in this way mediate a better theoretical understanding.
@@ 2.2 The time-causal limit kernel
over multiple temporal scale levels $\tau\in{\mathbb{R}}_{+}$ in such a way that no new local extrema or zero-crossings may be created from finer to coarser levels of temporal scales. The parameter $c$ is specifically a distribution parameter with $c>1$ that determines the discrete scale levels in units of the variance of the time-causal limit kernel according to [^2] This logarithmic spacing of the temporal scale levels is similar to the spacing generated by a uniform spacing with respect to the canonical logarithmic parameterization of the scale parameter into an effective scale parameter $\tau_{\text{eff}}$ according to $\tau_{\text{eff}}=A+B\log\tau$, as can be derived from axiomatic arguments for a regular spatial scale-space representation (Lindeberg 1993). Compared to a uniform spacing of the scale levels relative to the scale parameter measured in units of the variance $\tau$ of the temporal smoothing kernel, the logarithmic distribution according to (12) also leads to significantly shorter delays, as further described in Lindeberg (2016) Section 4 and Appendix 1. (Lindeberg 2016 Equation (18))
@@
where $\tau_{0}>0$ represents a reference scale, and where for common purposes of implementation one often chooses $\tau_{0}=1$ and either $c=\sqrt{2}$ or $c=2$. Alternatively, the time-causal limit kernel can also be parameterized in terms of the standard deviation $\sigma=\sqrt{\tau}$ of the time-causal limit kernel, which is of dimension $\mbox{time}]$.
@@ 2.4 Scale covariance property of the time-causal limit kernel
Temporal scale covariance here means that if we for any temporal scaling factor $S_{t}$, that is an integer power of the distribution parameter $c>1$ according to
$S_{t}=c^{j}\quad\quad\mbox{for}\quad\quad j\in{\mathbb{Z}},$ (20)
@@
This result thus means that the time-causal limit kernel is self-similar under temporal scaling transformations, which also, with appropriate complementary temporal scale normalization, carries over to temporal derivatives of the time-causal limit kernel according to
@@
Note, however, that since the temporal scale levels are inherently discrete, as implied by the construction of the temporal scale-space representation from multiple truncated exponential kernels coupled in cascade, the temporal scale covariance property holds exactly only when the temporal scaling factor $S_{t}$ is an integer power of the distribution parameter $c$ according to $S_{t}=c^{j}$ for $j\in{\mathbb{Z}}$. For other values of $S_{t}$, the corresponding result will only be approximate.
@@ 2.5.3 Trade-off issue regarding the choice of the distribution parameter $c$
  • A larger value of the distribution parameter $c$ will lead to shorter temporal delays, as estimated according to (28) and (29), with better ability to respond fast in real-time situations.
@@ 2.5.4 Differential equation formulation of the time-causal temporal scale-space representation
In terms of differential equations, [^4] The fact that the time-causal scale-space representation $L(t;\;\tau,c)$ according to (11) satisfies the recurrence relation (30) between adjacent levels of temporal scales can be formally verified as follows: Let us first rewrite the recurrence relation between the input signal $f_{\text{in}}(t)$ and the output signal $f_{\text{out}}(t)$ between adjacent levels of temporal scales $\tau_{k-1}$ and $\tau_{k}$ as $\partial_{t}f_{\text{out}}(t)=-1/\mu_{k}\times(f_{\text{out}}(t)-f_{\text{in}}(t))$. Taking the Laplace transform of this relation then gives $s\,{\cal L}(f_{\text{out}})(s)=-1/\mu_{k}\times({\cal L}(f_{\text{out}})(s)-{\cal L}(f_{\text{in}})(s))$. Let us next assume that the Laplace transforms ${\cal L}(f_{\text{out}})(s)$ and ${\cal L}(f_{\text{in}})(s)$ of $f_{\text{out}}(t)$ and $f_{\text{in}}(t)$, respectively, are related by a transfer function $H(s)$ according to ${\cal L}(f_{\text{out}})(s)=H(s)\,{\cal L}(f_{\text{in}})(s)$. Inserting this expression into the Laplace transform of the recurrence relation, then gives $s\,H(s)\,{\cal L}(f_{\text{in}})(s)=-1/\mu_{k}\times(H(s)\,{\cal L}(f_{\text{in}})(s)-{\cal L}(f_{\text{in}})(s))$, from which we can solve for the transfer function as $H(s)=1/(1+\mu_{k}\,s)$. From tables of the Laplace transforms of standard functions, we have that the Laplace transform of the truncated exponential kernel $e^{-\alpha\,t}$ for $t>0$, 0 elsewhere, is $1/(s+\alpha)$. From this, we can conclude that the transfer function $H(s)$ corresponds to the Laplace transform of the truncated exponential kernel $h_{\exp}(t;\;\mu_{k})=1/\mu_{k}\times e^{-t/\mu_{k}}$ for $t>0$, 0 elsewhere, which proves the result. the relationships between adjacent levels of temporal scales are given by (Lindeberg 2023 Equation (11))
@@
An important property of this type of temporal scale-space representation is that it is time-recursive. The temporal scale-space representations $L(t;\;\tau_{k},c)$ over the temporal scale channels with indices $k$ constitute a sufficient temporal memory of the past, to compute the temporal scale-space representation and the next temporal moment, given a new input in the input signal $f(t)$.
The form of the first-order integrators (30) also explicitly implies that local perturbations in the signal are gradually smoothed out from finer to coarser levels of temporal scales, because of the negative feedback with respect to the differences $L(t;\;\tau_{k},c)-L(t;\;\tau_{k-1},c)$ between the adjacent temporal scale channels.
@@ 2.5.5 Cascade smoothing property of the time-causal limit kernel
This property means that, when computing the temporal scale-space representation or a temporal derivative of the temporal scale-space representation, it is not necessary to redo all the computations when to compute those representations over multiple temporal scales. Instead, it is sufficient to perform the major amount of temporal smoothing once and for all, and then compute the representations at the coarser levels of temporal scales, by applying a usually much smaller number of truncated exponential kernels to the corresponding representations at finer temporal scales.
@@ 2.6 Time-causal and time-recursive mother wavelets and wavelet representations
According to traditional wavelet theory, one usually makes use of $L_{2}$-norms when normalizing the mother wavelet. For purposes of computing derivatives over multiple scales in scale-space theory, however, other choices of $L_{p}$-norms are more common for relating image structures over scales (Lindeberg 1998a Table 3), as will be detailed further below.
@@ 2.7 Relations to scale-normalized derivatives in scale-space theory
For purposes in scale-space theory, the notion of scale-normalized temporal derivatives can be introduced according to (Lindeberg 1998a Equation (6); Lindeberg 2017 Equation (6))
@@ 2.7.1 Scale covariance property of scale-normalized temporal derivatives
This means that we can readily compare the magnitudes of scale-normalized derivatives between matching temporal moments $\{t\leftrightarrow t^{\prime}\}$ and temporal scales $\{\tau\leftrightarrow\tau^{\prime}\}$.
@@ 2.7.2 Scale selection properties for characteristic temporal signals
In this way, the scale selection mechanism based on local extrema over temporal scales leads to scale estimates $\hat{\tau}$ that reflect the inherent temporal scales $\tau$ in the input signal. Notably, these results bear close relationships to the matched filter theorem (Woodward 1953, Turin 1960), in that the scale selection mechanism will choose filters for detecting the different types of temporal structures in the input signal, that as well as possible match their size.
@@
In relation to normalization of the filters with respect to $L_{p}$-norms, it can specifically be shown that (Lindeberg 1998a Section 9.1) the notion of $\gamma$-normalized derivatives corresponds to normalizing the $n$:th order scale-normalized Gaussian derivative kernels $g_{\zeta^{n}}(t;\;\tau)$ over a one-dimensional domain to constant $L_{p}$-norms over scale
@@ 3.1 Time-causal wavelet representations based on temporal derivatives of the time-causal limit kernel
for the temporal scale levels chosen as (Lindeberg 2016 Equation (18))
@@
In the following analysis to be performed, we will, however, not explicitly normalize the mother wavelets to constant $L_{p}$-norm, instead fully relying on the up to a uniform scaling factor normalization over scales in terms of scale-normalized temporal derivatives according to (45).